# Thomason cohomology and Quillen’s Theorem A

### v1

@@ 1. Introduction and statement of results
$\kappa:\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)\to N\mathcal{C}.$
@@
For a small category $\mathcal{C}$, let $N\mathcal{C}$ denote the nerve of the category $\mathcal{C}$. We can consider the simplex category $\Delta(N\mathcal{C})$ of the simplicial set $N\mathcal{C}$. A functor $\Delta(N\mathcal{C})\to R$-Mod is called a coefficient system on $N\mathcal{C}$ (see Section 3.2 for details). The Thomason cohomology $H^{*}_{Th}(\mathcal{C};\mathcal{M})$ of a category $\mathcal{C}$ with coefficients in $\mathcal{M}$ is defined as the cohomology of the simplicial set $N\mathcal{C}$ with coefficients in $\mathcal{M}$ which is isomorphic to the cohomology of the category of simplices $\Delta(N\mathcal{C})$ of the nerve $N\mathcal{C}$ with coefficients in $\mathcal{M}$ (see Section 3.4). The cohomology groups $H^{*}(\mathcal{C};M)$ and $H^{*}_{BW}(\mathcal{C};M)$ can be realized as the Thomason cohomology $H^{*}_{Th}(\mathcal{C};\mathcal{M})$ of $\mathcal{C}$ with suitable choices of coefficient systems $\mathcal{M}$ (see Lemma 3.7).
@@
[…]
@@
where $\mathcal{M}_{d}=(j_{d}^{\prime})^{*}\mathcal{M}$. Then the induced map
@@
Let $F:\mathcal{D}\to Cat$ be a functor and $\int_{\mathcal{D}}F$ be the Grothendieck construction of $F$. Let $\pi:\int_{\mathcal{D}}F\to\mathcal{D}$ be the canonical functor that sends a pair $(d,x)\in\int_{\mathcal{D}}F$ to $d\in\mathcal{D}$, and $\widetilde{F}:\mathcal{D}\to Cat$ be the functor that sends $d\in\mathcal{D}$ to the comma category $\pi/d$. Then for every coefficient system $\mathcal{M}$ for $N\mathcal{C}$, there is a spectral sequence
@@
Acknowledgements: The first author is supported by TÜBİTAK grant 2211-A National PhD Scholarship Program. The second author is supported by TÜBİTAK grant 2219-International Postdoctoral Research Fellowship Program (2023, 2nd term). We thank TÜBİTAK for their support of this research. We also thank Matthew Gelvin and Caroline Yalçın for reading an earlier version of the paper and for their suggestions.
@@ 2. Modules and cohomology for small categories
This shows that the $R\mathcal{C}$-module $R\mathrm{Mor}_{\mathcal{C}}(x,?)$ is projective (see [14], [20] for details). Having enough projective modules allows us to construct projective resolutions and define ext-groups over $R\mathcal{C}$-modules.
@@
Instead of covariant functors $M:\mathcal{C}\to R$-Mod, one can consider the module category of contravariant functors $M:\mathcal{C}^{op}\to R$-Mod. We call such a module a contravariant $R\mathcal{C}$-module (or a right $R\mathcal{C}$-module). The contravariant $R\mathcal{C}$-modules defined by $R\mathrm{Mor}_{\mathcal{C}}(?,x)$ for $x\in\mathcal{C}$ are projective $R\mathcal{C}$-modules, and the ext-groups $\mathrm{Ext}^{*}_{R\mathcal{C}}(M,N)$ for contravariant $R\mathcal{C}$-modules $M$ and $N$ are defined in a similar way using contravariant projective $R\mathcal{C}$-modules.
@@ 2.1. Cohomology of small categories
$(\delta^{n-1}f)(\sigma)=\sum_{i=0}^{n}(-1)^{i}f(d_{i}\sigma)+(-1)^{n}M(\alpha_{n})f(d_{n}\sigma)$
@@ 2.2. Baues-Wirsching Cohomology
${x\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{u}$${x^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{v}$$\scriptstyle{\alpha^{\prime}}$${y^{\prime}}$
@@
Let $\mathcal{C}$ be a small category and $M$ be a natural system for $\mathcal{C}$ over $R$ The Baues-Wirsching cohomology $H^{*}_{BW}(\mathcal{C};M)$ of $\mathcal{C}$ with coefficients in $M$ is the cohomology of a cochain complex $C^{*}_{BW}(\mathcal{C};M)$ where for all $n\geq 0$, $C^{n}_{BW}(\mathcal{C};M)$ is the abelian group of all functions
@@
$\delta^{n}(f)(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})$ $=\alpha_{1}^{*}f(c_{1}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{2}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})$
@@
To prove Proposition 2.7, Baues and Wirsching construct a projective resolution $P_{*}\to\underline{R}$ of the constant functor $\underline{R}$ as a $R\mathfrak{F}\mathcal{C}$-module such that $\mathrm{Hom}_{RF\mathcal{C}}(P_{*},M)$ is a cochain complex isomorphic to $C^{*}_{BW}(\mathcal{C};M)$. Note that for every object $\alpha\in\mathfrak{F}\mathcal{C}$, the $R\mathfrak{F}\mathcal{C}$-module $R\mathrm{Mor}_{\mathfrak{F}\mathcal{C}}(\alpha,?)$ is a projective $R\mathfrak{F}\mathcal{C}$-module. One can use these modules to construct the projective resolution (see [1, Theorem 4.4] for details).
@@
[…]
@@
[…]
@@ 3.1. Coefficient systems for simplicial sets
The simplex category $\Delta$ is the category whose objects are ordered finite sets $[n]=\{0,1,\dots,n\}$, and whose morphisms are given by weakly order-preserving functions between them. Morphisms in the simplex category are generated by coboundary maps $d^{i}:[n-1]\to[n]$, $0\leq i\leq n$, and codegeneracy maps $s^{i}:[n+1]\to[n]$, $0\leq i\leq n$, which satisfy certain relations (see [9]).
@@ 3.2. Cohomology of simplicial sets with general coefficients
for every $f\in C^{n-1}(X;\mathcal{M})$. The cohomology $H^{*}(X;\mathcal{M})$ of the simplicial set $X$ with coefficients in $\mathcal{M}$ is the cohomology of the cochain complex $C^{*}(X;M)$.
@@ 3.3. Induced Maps on Cohomology
[…]
@@ 3.4. The Thomason cohomology of a category
[…]
@@ 4.1. Homotopy colimits
$\mathrm{diag}X:\Delta^{op}\to\Delta^{op}\times\Delta^{op}\to Sets$
where the first functor is the diagonal functor defined by $[n]\to([n],[n])$.
@@
$(f_{h},f_{v})^{*}(\sigma,\tau)=(f_{h}^{*}(\sigma),F(\alpha_{f_{h}(0)}\cdots\alpha_{1})(f_{v}^{*}(\tau))).$
@@ 4.2. The Dold-Puppe Theorem
The Moore complex of a simplicial $R$-module $A$ is the chain complex whose $n$-chains are given by $A_{n}$ and whose boundary maps $\partial_{n}:A_{n}\to A_{n-1}$ are defined by $\partial_{n}(x)=\sum_{i=0}^{n}(-1)^{i}d_{i}(x)$. The Moore complex of $A$, denoted by $A_{*}$, is a chain complex and its homology is denoted by $H_{*}(A)$.
@@
$\partial^{h}_{p}=\sum_{i=0}^{p}(-1)^{i}d_{i}^{h}\ \text{ and }\ \partial^{v}_{q}=\sum_{j=1}^{q}(-1)^{j}d_{j}^{v}.$
@@
$\partial_{n}(x)=\partial^{h}_{p}(x)+(-1)^{q}\partial_{q}^{v}(x)$
@@
Another way to get a chain complex from a bisimplicial $R$-module $A$ is to consider the simplicial $R$-module $\mathrm{diag}A$ and take its Moore complex $(\mathrm{diag}A)_{*}$. By the Dold-Puppe theorem, these two different constructions give chain complexes that are homotopic.
@@
Let $\varphi:\mathcal{C}\to\mathcal{D}$ be a functor between two small categories, and $R$ be a commutative ring. We denote by $G:=(\varphi/-):\mathcal{D}\to Cat$ the functor that sends $d\in\mathcal{D}$ to the comma category $\varphi/d$. Consider the simplicial replacement $N(\mathcal{D};G)$ of $G$. For each $d\in\mathcal{D}$, we can identify $NG(d)_{q}$ with the set of pairs $(\tau,\mu)$ where $\tau=(c_{0}\to\cdots\to c_{q})\in N\mathcal{C}_{q}$ and $\mu:\varphi(c_{q})\to d$ is a morphism in $\mathcal{D}$. This identification allows us to write the $(p,q)$-simplices in $N(\mathcal{D};G)$ as the triples $(\sigma,\tau,\mu)$ where
@@
The homotopy colimit
$X:=\operatorname*{hocolim}_{\mathcal{D}}G$
is the diagonal of the bisimplicial set $N(\mathcal{D};G)$. Let $\kappa:X\to N\mathcal{C}$ be the simplicial map defined by $\kappa(\sigma,\tau,\mu)=\tau$. Our main aim in this section is to prove Theorem 1.2: For any coefficient system $\mathcal{M}:\Delta(N\mathcal{C})\to R$-Mod, the induced map
@@
$X_{p,q}(\theta)=\{(\sigma,\tau,\mu,f)\,|\,(\sigma,\tau,\mu)\in N(\mathcal{D};G)_{p.q},f:[q]\to[n],f^{*}(\theta)=\tau\}.$
@@
$\mathrm{Hom}_{R\Delta(N\mathcal{C})}((\mathrm{diag}\mathbb{P})_{k},\mathcal{M})$ $\cong\prod_{(\sigma,\tau,\mu)\in N(\mathcal{D};G)_{k,k}}\mathrm{Hom}_{R\Delta(N\mathcal{C})}(R\mathrm{Mor}_{\Delta(N\mathcal{C})}(\tau,?),\mathcal{M})$
$\cong\prod_{(\sigma,\tau,\mu)\in N(\mathcal{D};G)_{k,k}}\mathcal{M}(\tau)$
$\cong C^{k}(\mathrm{diag}N(\mathcal{D};G);\kappa^{*}(\mathcal{M})).$
@@
$\mathrm{Hom}_{R\Delta(N\mathcal{C})}((\mathrm{diag}\mathbb{Q})_{k},\mathcal{M})$ $\cong\prod_{\tau\in N\mathcal{C}_{q}}\mathrm{Hom}_{R\Delta(N\mathcal{C})}(R\mathrm{Mor}_{\Delta(N\mathcal{C})}(\tau,?),\mathcal{M})$
@@
[…]
where the first vertical map is induced by the bisimplicial map $\xi:\mathbb{P}\to\mathbb{Q}$ and the second vertical map is induced by the simplicial map $\kappa:N(\mathcal{D};G)\to N\mathcal{C}$. Since the augmented complexes for $(\mathrm{diag}\mathbb{P})_{*}$ and $(\mathrm{diag}\mathbb{Q})_{*}$ are projective resolutions for the constant functor $\underline{R}$, the first vertical map $\xi^{*}$ induces an isomorphism on cohomology. Hence $\kappa^{*}$ also induces an isomorphism on cohomology. ∎
@@
Let $\varphi:\mathcal{C}\rightarrow\mathcal{D}$ be a functor between two small categories and \$-\backslash\varphi:\mathcal{D}^{op}\rightarrow Cat$ denote the functor that sends each $d\in\mathcal{D}$ to the comma category $d\backslash\varphi$. Then for every coefficient system $\mathcal{M}:\Delta(N\mathcal{C})\rightarrow R$-Mod, the homotopy equivalence $\upsilon:Y=\operatorname*{hocolim}_{\mathcal{D}^{op}}N(-\backslash\varphi)\to N\mathcal{C}$ induces an isomorphism

### v2

@@ Abstract.
Corresponding Author: Ergün Yalçın, Email: yalcine@fen.bilkent.edu.tr
Address: Bilkent University, Department of Mathematics, 06800, Bilkent, Ankara, Türkiye
Contributing Author: Mehmet Kırtışoğlu, Email: m.kirtisoglu@bilkent.edu.tr
Address: Bilkent University, Department of Mathematics, 06800, Bilkent, Ankara, Türkiye
@@ 1. Introduction and statement of results
$\kappa:\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)\to N\mathcal{C}$
defined by $\kappa(\sigma,\tau,\mu)=\tau$ (see Lemma 5.1).
@@
For a small category $\mathcal{C}$, let $N\mathcal{C}$ denote the nerve of $\mathcal{C}$. A functor $\mathcal{M}:\Delta(N\mathcal{C})\to R$-Mod is called a coefficient system for $N\mathcal{C}$ where $\Delta(N\mathcal{C})$ denotes the category of simplices of $N\mathcal{C}$ (see Definition 3.1). For a coefficient system $\mathcal{M}$ for $N\mathcal{C}$, the Thomason cohomology $H^{*}_{Th}(\mathcal{C};\mathcal{M})$ is defined as the Quillen cohomology of the category $\Delta(N\mathcal{C})$ with coefficients in $\mathcal{M}$ (see Section 3.4). The cohomology groups $H^{*}(\mathcal{C};M)$ and $H^{*}_{BW}(\mathcal{C};M)$ can be realized as the Thomason cohomology $H^{*}_{Th}(\mathcal{C};\mathcal{M})$ of $\mathcal{C}$ with suitable choices of coefficient systems $\mathcal{M}$ (see Lemma 3.7).
@@
${N(\varphi/d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{d}}$$\scriptstyle{NU(d)}$${N\mathcal{C}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{N\varphi}$${N(\mathrm{id}/d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{d}^{\prime}}$${N\mathcal{D}}$
@@
where $\mathcal{M}_{d}=(j_{d}^{\prime})^{*}\mathcal{M}$. Then the map
@@
Let $F:\mathcal{D}\to Cat$ be a functor and $\int_{\mathcal{D}}F$ be the Grothendieck construction of $F$. Let $\pi:\int_{\mathcal{D}}F\to\mathcal{D}$ be the canonical functor that sends a pair $(d,x)\in\int_{\mathcal{D}}F$ to $d\in\mathcal{D}$, and $\widetilde{F}:\mathcal{D}\to Cat$ be the functor that sends $d\in\mathcal{D}$ to the comma category $\pi/d$. Then for every coefficient system $\mathcal{M}$ defined on $N(\int_{\mathcal{D}}F)$, there is a spectral sequence
@@ 2. Modules and cohomology for small categories
Using this isomorphism, one can show that the $R\mathcal{C}$-module $R\mathrm{Mor}_{\mathcal{C}}(x,?)$ is projective (see [14], [20] for details). Having enough projective modules allows us to construct projective resolutions and define ext-groups over $R\mathcal{C}$-modules.
@@
Instead of covariant functors $M:\mathcal{C}\to R$-Mod, one can consider the abelian category of contravariant functors $M:\mathcal{C}^{op}\to R$-Mod. We call such a module a contravariant $R\mathcal{C}$-module (or a right $R\mathcal{C}$-module). The contravariant $R\mathcal{C}$-modules defined by $R\mathrm{Mor}_{\mathcal{C}}(?,x)$ for $x\in\mathcal{C}$ are projective, and the ext-groups $\mathrm{Ext}^{*}_{R\mathcal{C}}(M,N)$ for contravariant $R\mathcal{C}$-modules $M$ and $N$ are defined in a similar way using the resolutions of contravariant $R\mathcal{C}$-modules.
@@ 2.1. Cohomology of small categories
$(\delta^{n-1}f)(\sigma)=\sum_{i=0}^{n-1}(-1)^{i}f(d_{i}\sigma)+(-1)^{n}M(\alpha_{n})f(d_{n}\sigma)$
@@ 2.2. Baues-Wirsching Cohomology
[…]
@@
Let $\mathcal{C}$ be a small category and $M$ be a natural system for $\mathcal{C}$ over $R$ . The Baues-Wirsching cohomology $H^{*}_{BW}(\mathcal{C};M)$ of $\mathcal{C}$ with coefficients in $M$ is the cohomology of a cochain complex $C^{*}_{BW}(\mathcal{C};M)$ where for all $n\geq 0$, $C^{n}_{BW}(\mathcal{C};M)$ is the abelian group of all functions
@@
$\delta^{n-1}(f)(c_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})$ $=\alpha_{1}^{*}f(c_{1}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{2}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{n}}\,}c_{n})$
@@
To prove Proposition 2.7, Baues and Wirsching construct a projective resolution $P_{*}\to\underline{R}$ of the constant functor $\underline{R}$ as a $R\mathfrak{F}\mathcal{C}$-module such that $\mathrm{Hom}_{R\mathfrak{F}\mathcal{C}}(P_{*},M)$ is a cochain complex isomorphic to $C^{*}_{BW}(\mathcal{C};M)$. Note that for every object $\alpha\in\mathfrak{F}\mathcal{C}$, the $R\mathfrak{F}\mathcal{C}$-module $R\mathrm{Mor}_{\mathfrak{F}\mathcal{C}}(\alpha,?)$ is projective. One can use these modules to construct the projective resolution (see [1, Theorem 4.4] for details).
@@
$M:\mathfrak{F}\mathcal{C}\smash{\,\mathop{\longrightarrow}\limits^{T}\,}\mathcal{C}\smash{\,\mathop{\longrightarrow}\limits^{N}\,}R\text{-mod},$
@@
$M:\mathfrak{F}\mathcal{C}\smash{\,\mathop{\longrightarrow}\limits^{S}\,}\mathcal{C}^{op}\smash{\,\mathop{\longrightarrow}\limits^{N^{\prime}}\,}R\text{-mod},$
@@ 3.1. Coefficient systems for simplicial sets
The simplex category $\Delta$ is the category whose objects are ordered finite sets $n]=\{0,1,\dots,n\}$, and whose morphisms are given by weakly order-preserving functions between them. Morphisms in the simplex category are generated by coboundary maps $d^{i}:[n-1]\to[n]$, $0\leq i\leq n$, and codegeneracy maps $s^{i}:[n+1]\to[n]$, $0\leq i\leq n$, which satisfy certain relations (see [9]).
@@ 3.2. Cohomology of simplicial sets with general coefficients
for every $f\in C^{n-1}(X;\mathcal{M})$ and $\sigma\in X_{n}$. The cohomology $H^{*}(X;\mathcal{M})$ of the simplicial set $X$ with coefficients in $\mathcal{M}$ is the cohomology of the cochain complex $C^{*}(X;M)$.
@@ 3.3. Induced Maps on Cohomology
$\lambda^{*}\mathcal{M}:\Delta(X)\xrightarrow{\Delta(\lambda)}\Delta(Y)\xrightarrow{\mathcal{M}}R\text{-Mod}$
@@ 3.4. The Thomason cohomology of a category
$\mathcal{M}:\Delta(N\mathcal{C})\smash{\,\mathop{\longrightarrow}\limits^{\chi}\,}\mathfrak{F}\mathcal{C}\smash{\,\mathop{\longrightarrow}\limits^{M}\,}R\text{-Mod}$
@@ 4.1. Homotopy colimits
$\mathrm{diag}X:\Delta^{op}\to\Delta^{op}\times\Delta^{op}\to\mathcal{C}$
where the first functor is the diagonal functor defined by $n]\to([n],[n])$.
@@
$(f_{h},f_{v})^{*}(\sigma,\tau)=(f_{h}^{*}(\sigma),F(\alpha_{f_{h}(0)}\cdots\alpha_{1})(f_{v}^{*}(\tau))).$ (1)
@@ 4.2. The Dold-Puppe Theorem
The Moore complex of a simplicial $R$-module AA is the chain complex whose $n$-chains are given by $A_{n}$ and whose boundary maps $\partial_{n}:A_{n}\to A_{n-1}$ are defined by $\partial_{n}(x)=\sum_{i=0}^{n}(-1)^{i}d_{i}(x)$. The Moore complex of $A$, denoted by $A_{*}$, is a chain complex and its homology is denoted by $H_{*}(A)$.
@@
$\partial^{h}_{p}=\sum_{i=0}^{p}(-1)^{i}d_{i}^{h}\ \text{ and }\ \partial^{v}_{q}=\sum_{j=0}^{q}(-1)^{j}d_{j}^{v}.$
@@
$\partial_{n}(x)=\partial^{h}_{p}(x)+(-1)^{p}\partial_{q}^{v}(x)$
@@
Another way to get a chain complex from a bisimplicial $R$-module AA is to consider the simplicial $R$-module $\mathrm{diag}A$ and take its Moore complex $(\mathrm{diag}A)_{*}$. By the Dold-Puppe theorem, these two different constructions give chain complexes that are homotopic.
@@
Let $\varphi:\mathcal{C}\to\mathcal{D}$ be a functor between two small categories, and $R$ be a commutative ring. We denote by $G:=(\varphi/-):\mathcal{D}\to Cat$ the functor that sends $d\in\mathcal{D}$ to the comma category $\varphi/d$. Let $NG:\mathcal{D}\to sSet$ denote the functor that sends $d\in\mathcal{D}$ to the nerve $NG(d)$. Consider the simplicial replacement $N(\mathcal{D};NG)$ of $NG$. For each $d\in\mathcal{D}$, we can identify $NG(d)_{q}$ with the set of pairs $(\tau,\mu)$ where $\tau=(c_{0}\to\cdots\to c_{q})\in N\mathcal{C}_{q}$ and $\mu:\varphi(c_{q})\to d$ is a morphism in $\mathcal{D}$. This identification allows us to write the $(p,q)$-simplices in $N(\mathcal{D};NG)$ as the triples $(\sigma,\tau,\mu)$ where
@@
Lemma 5.1.
Consider $N\mathcal{C}$ as a bisimplicial set by taking $(N\mathcal{C})_{p,q}=N\mathcal{C}_{q}$ for all $p$, and by taking all the horizontal maps as the identity maps. Then the map $\kappa:N(\mathcal{D};NG)\to N\mathcal{C}$ defined by $\kappa(\sigma,\tau,\mu)=\tau$ is a map of bisimplicial sets.
Proof.
This follows from the formula
$(f_{h}\times f_{v})^{*}(\sigma,\tau,\mu)=(f_{h}^{*}(\sigma),f_{v}^{*}(\tau),\mu^{\prime})$
given above in (3). ∎
Let $X:=\operatorname*{hocolim}_{\mathcal{D}}NG$ denote the homotopy colimit of the functor $NG$, and let $\kappa:X\to N\mathcal{C}$ be the map defined by $\kappa(\sigma,\tau,\mu)=\tau$. By Lemma 5.1, $\kappa$ is a simplicial map obtained as a restriction of a bisimplicial map. Our main aim in this section is to prove Theorem 1.2 which states that for any coefficient system $\mathcal{M}:\Delta(N\mathcal{C})\to R$-Mod, the induced map
@@
$X_{p,q}(\theta)=\{(\sigma,\tau,\mu,f)\,|\,(\sigma,\tau,\mu)\in N(\mathcal{D};NG)_{p.q},f:[q]\to[n],f^{*}(\theta)=\tau\}.$
@@
$\mathrm{Hom}_{R\Delta(N\mathcal{C})}((\mathrm{diag}\mathbb{P})_{k},\mathcal{M})$ $\cong\prod_{(\sigma,\tau,\mu)\in N(\mathcal{D};NG)_{k,k}}\mathrm{Hom}_{R\Delta(N\mathcal{C})}(R\mathrm{Mor}_{\Delta(N\mathcal{C})}(\tau,?),\mathcal{M})$
$\cong\prod_{(\sigma,\tau,\mu)\in N(\mathcal{D};NG)_{k,k}}\mathcal{M}(\tau)$
$\cong C^{k}(\mathrm{diag}N(\mathcal{D};NG);\kappa^{*}(\mathcal{M})).$
@@
$\mathrm{Hom}_{R\Delta(N\mathcal{C})}((\mathrm{diag}\mathbb{Q})_{k},\mathcal{M})$ $\cong\prod_{\tau\in N\mathcal{C}_{k}}\mathrm{Hom}_{R\Delta(N\mathcal{C})}(R\mathrm{Mor}_{\Delta(N\mathcal{C})}(\tau,?),\mathcal{M})$
@@
${\mathrm{Hom}_{R\Delta(N\mathcal{C})}((\mathrm{diag}\mathbb{P})_{*},\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${C^{*}(\mathrm{diag}N(\mathcal{D};NG);\kappa^{*}(\mathcal{M}))}$${\mathrm{Hom}_{R\Delta(N\mathcal{C})}((\mathrm{diag}\mathbb{Q})_{*},\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$$\scriptstyle{\xi^{*}}$${C^{*}(N\mathcal{C};\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\kappa^{*}}$
where the first vertical map is induced by the bisimplicial map $\xi:\mathbb{P}\to\mathbb{Q}$ and the second vertical map is induced by the simplicial map $\kappa:N(\mathcal{D};NG)\to N\mathcal{C}$. Since the augmented complexes for $(\mathrm{diag}\mathbb{P})_{*}$ and $(\mathrm{diag}\mathbb{Q})_{*}$ are projective resolutions for the constant functor $\underline{R}$, the first vertical map $\xi^{*}$ induces an isomorphism on cohomology. Hence $\kappa^{*}$ also induces an isomorphism on cohomology. ∎
@@
Let $\varphi:\mathcal{C}\rightarrow\mathcal{D}$ be a functor between two small categories and $\backslash\varphi:\mathcal{D}^{op}\rightarrow Cat$ denote the functor that sends each $d\in\mathcal{D}$ to the comma category $d\backslash\varphi$. Then for every coefficient system $\mathcal{M}:\Delta(N\mathcal{C})\rightarrow R$-Mod, the homotopy equivalence $\upsilon:Y=\operatorname*{hocolim}_{\mathcal{D}^{op}}N(-\backslash\varphi)\to N\mathcal{C}$ induces an isomorphism